# MM-Soc: Benchmarking Multimodal Large Language Models in Social Media Platforms

### v1

@@ 7 Conclusion
Our study presents a comprehensive evaluation of $4$ leading MLLMs on $10$ carefully constructed multimodal social media tasks from diverse domains such as misinformation, hate speech, memes, and a novel YouTube dataset, which comprises our proposed MM-Soc benchmark. Our evaluation of the current capabilities presents the following insights: (i) zero-shot capabilities of certain MLLMs are near-random and underperform drastically in comparison to smaller fully fine-tuned models, (ii) LLaVA-v1.5 is the most competitive open-source MLLM so far, and (iii) instruction following capabilities of MLLMs improve with their size. MM-Soc also enables quantitative case studies, two of which were illustrated in this work and revealed (a) the limitations of MLLMs in self-improving their accuracy and (b) the effectiveness of fine-tuning MLLMs with labeled data. As benchmarks highlight current limitations and guide future research, we intend to expand MM-Soc’s coverage to more models and social media tasks to encourage reliable applicability of MLLMs in online spheres.
@@ 8 Limitations
We address some limitations of the current study settings, while discussing potential directions for future works.
@@ 8.1 Exclusion of Proprietary Models
We excluded models like GPT4V and Gemini from our study for specific reasons. First, this research aims to spotlight the constraints of open-source MLLMs in tackling multimodal tasks derived from social media contexts. This emphasis on open-source models is driven by our commitment to enhancing privacy protection. Unlike proprietary models that aggregate data of multiple platforms onto a central server, posing significant privacy risks and operational costs, open-source models are able to process data in a decentralized way Fan et al. (2023); Zhang et al. (2023c). This distinction not only ensures better privacy safeguards but also resonates with our objective to spotlight and scrutinize the limitations inherent within open-source frameworks when deployed in complex, real-world scenarios like social media. By doing so, we hope that the research community can dedicate resources towards the development of more sophisticated open-source models that address these gaps, promoting the ethos of open science. Second, proprietary models like Gemini reject images containing people and prompts associated with misinformation and hate speech. These restrictions present significant barriers to a comprehensive analysis of MLLMs’ performance in handling the diverse and often complex content found on social media platforms.
@@ 8.2 Scope of Datasets Included in Benchmark
9 Ethical Considerations & Broader Impacts
@@
Appendix A Details on Datasets
A.1 Tagging
The tagging task focuses on predicting appropriate “topic categories” for YouTube videos, chosen from a predefined set as listed in Table 7. These topics make it easier for users to find videos that match their interests but also enhance the overall content management strategy Jin et al. (2023b). This dataset exemplifies the necessity of multimodal understanding in categorizing online video content. The dataset is licensed under the Apache 2.0 License.
A.2 Misinformation datasets
@@ A.4 Memotion
Appendix B Details on Experiments
B.1 Evaluation Metrics
@@
Model Pre Rec F1 Jaccard $\mathcal{L}_{\mathrm{Hamming}}$downarrow
instructblip-flan-t5-xl 0.045 0.326 0.057 0.036 0.500
instructblip-flan-t5-xxl 0.092 0.376 0.128 0.078 0.161
instructblip-vicuna-13b 0.044 0.230 0.050 0.032 0.429
blip2-opt-2.7b 0.027 0.037 0.022 0.013 0.223
blip2-flan-t5-xl 0.196 0.191 0.157 0.112 0.092
blip2-flan-t5-xxl 0.176 0.350 0.188 0.122 0.085
llama-adapter-v2 0.028 0.029 0.021 0.012 0.137
llava-v1.5-7b 0.048 0.345 0.068 0.041 0.406
 + finetuning on ground-truth 0.162 0.373 0.209 0.148 0.063
 + finetuning on explanations 0.562 0.491 0.494 0.400 0.027
llava-v1.5-13b 0.052 0.361 0.071 0.043 0.342
 + finetuning on ground-truth 0.123 0.441 0.167 0.113 0.104
 + finetuning on explanations 0.533 0.473 0.474 0.387 0.027
Table 8: Results of tagging on the YouTube dataset. “FT-Labels” and “FT-Explanations” represent the models fine-tuned on the ground-truth labels and explanations, respectively. A $\downarrow$” next to the metric indicates that lower values represent better performances. instructblip-vicuna-7b fails to produce valid predictions in this context.
@@ B.2 Details on Models
Table 11: Accuracy of all models on the tasks. Best and 2nd best performances among the MLLMs are highlighted in bold and underline, respectively. “ID” and “SCD” stand for the image description task and the social context description task, respectively.
B.3 Implementation Details
Benchmark Evaluation For inference, we use Nucleus Sampling Holtzman et al. (2019) with a probability threshold of 0.9, a temperature of 1.0, and a maximum output length of 256 tokens. To account for the randomness in the generation process, we run each experiment with 3 random seeds and report the average results. All experiments were conducted on a server with 5 A100 80GB GPUs.
The models are implemented using the Transformers library Wolf et al. (2020). We use the NLTK package Loper and Bird (2002) to calculate BLEU scores, the rouge [^4] https://github.com/pltrdy/rouge package to calculate ROUGE scores and the sentence-bert [^5] https://github.com/UKPLab/sentence-transformers package to calculate sentence embedding similarities, respectively.

### v2

@@ 6 Conclusion
Our study presents a comprehensive evaluation of $4$ leading MLLMs on $10$ carefully constructed multimodal social media tasks from diverse domains such as misinformation, hate speech, memes, and a novel YouTube dataset, which comprises our proposed MM-Soc benchmark. Our evaluation of the current capabilities presents the following insights: (i) zero-shot capabilities of certain MLLMs are near-random and underperform drastically in comparison to smaller fully fine-tuned models, (ii) LLaVA-v1.5 is currently the most competitive open-source MLLM, and (iii) instruction following capabilities of MLLMs improve with their size. MM-Soc also enables quantitative case studies, two of which were illustrated in this work and revealed (a) the limitations of MLLMs in self-improving their accuracy and (b) the effectiveness of fine-tuning MLLMs with labeled data. As benchmarks highlight current limitations and guide future research, we intend to expand MM-Soc’s coverage to more models and social media tasks to encourage reliable applicability of MLLMs in online spheres.
@@ 7 Limitations
We describe limitations of the current study settings and discuss potential directions for future works.
@@ 7.1 Exclusion of Proprietary Models
This study does not focus on proprietary models like GPT-4V and Gemini for specific reasons. First, this research aims to spotlight the constraints of open-source MLLMs in tackling multimodal tasks derived from social media contexts. This emphasis on open-source models is driven by our commitment to enhancing privacy protection. Unlike proprietary models that aggregate data of multiple platforms onto a central server, posing significant privacy risks and operational costs, open-source models are able to process data in a decentralized way Fan et al. (2023); Zhang et al. (2023c). This distinction not only ensures better privacy safeguards but also resonates with our objective to spotlight and scrutinize the limitations inherent within open-source frameworks when deployed in complex, real-world scenarios like social media. By doing so, we hope that the research community can dedicate resources towards the development of more sophisticated open-source models that address these gaps, promoting the ethos of open science. Second, proprietary models like Gemini reject images containing people and prompts associated with misinformation and hate speech. These restrictions present significant barriers to a comprehensive analysis of MLLMs’ performance in handling the diverse and often complex content found on social media platforms.
@@ 7.2 Scope of Datasets Included in Benchmark
8 Ethical Considerations and Broader Impacts
@@
Acknowledgements
This research/material is based upon work supported in part by NSF grants CNS-2154118, ITE-2137724, ITE-2230692, CNS2239879, Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112290102 (subcontract No. PO70745), CDC, and funding from Microsoft. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the position or policy of DARPA, DoD, SRI International, CDC, NSF, and no official endorsement should be inferred. We thank members of the CLAWS Lab for their helpful feedback.
Appendix A Task Selection
The selection of tasks and datasets in MM-Soc centers around three key criteria:
  • Tasks that require multimodal understanding of both textual and image domains;
  • Tasks directly related to the dynamics of social media platforms;
  • Tasks that have undergone rigorous evaluation in subsequent research, which affirms their validity as a benchmark.
The task selection process started with a comprehensive literature review through NLP conferences (ACL, EMNLP, NAACL, SemEval), Machine Learning conferences (NeurIPS and ICML) and Data Mining conferences (KDD and SIGIR) since 2019. Papers satisfying these criteria were retained. Our final list of tasks, while collectively categorized under multimodal engagements in social media contexts, each distinctly require a variety of cognitive capabilities. Some of these capabilities intersect across different tasks, while others are unique to specific challenges. Every task demands that models not only comprehend textual instructions but also accurately interpret relevant visual information to solve the task.
Appendix B The YouTube2M dataset
B.1 Distribution of Languages
Table 8 shows the language distribution of YouTube videos in YouTube2M. There are 138 unique languages in YouTube2M. 323,007 videos have explicitly specified their default languages, representing 16.45% of the total 1,963,697 videos. We provide a detailed breakdown of the languages, showcasing the distribution of the top 10 most and least popular languages within the dataset. Our findings reveal a long-tail distribution in language popularity. Notably, English (including en, en-GB, en-US) dominates the dataset with 275,408 videos, accounting for 85.3% of videos with a specified language. In contrast, the ten least common languages each only appear once.
Tag Count
Music 524,369
Video_game_culture 475,124
Action_game 406,582
Lifestyle_(sociology) 383,092
Action-adventure_game 344,743
Role-playing_video_game 320,837
Entertainment 201,722
Strategy_video_game 200,968
Society 190,023
Pop_music 171,038
Rock_music 158,922
Table 9: The most popular tags in YouTube2M
B.2 Distribution of Tags
The YouTube2M dataset encompasses a rich variety of 62 unique tags, with 1,389,219 videos bearing the top 5 tags, as shown in Table 9. Note that a video can have multiple tags. This accounts for 70.7% of the entire dataset. We observe a strong inclination towards gaming and music content.
B.3 Channel Information
Channel Description #Videos
euronews European television news network, headquartered in Brussels, Belgium 4737
Al Jazeera English 24-hour English news channel headquartered in Doha, Qatar 4655
IGN American video game media featuring videos about the latest gaming & entertainment news and events 4041
Fox News 24-hour all-encompassing news service dedicated to delivering breaking news and political & business news 3504
WWE World Wrestling Entertainment (WWE) original shows 2847
The Young Turks American progressive news commentary show 2345
ESPN Multimedia sports entertainment channel 2026
Movieclips Online movie clips collection 1883
CNN Multinational news channel and website 1808
The Majority Report w/ Sam Seder daily political talk show 1655
Table 10: YouTube channels with the most videos in the dataset.
There are 604,340 unique channels associated with the videos in the dataset. The most popular 10 channels and their associated videos are shown in Table 9. As observed from the statistics, a large portion of videos propagated in online social networks are centered around news and sports, signifying the popularity of these topics among online discourse.
Appendix C Details about Datasets
C.1 Tagging
The tagging task focuses on predicting appropriate “topic categories” for YouTube videos, chosen from a predefined set as listed in Table 7. These topics make it easier for users to find videos that match their interests but also enhance the overall content management strategy. This dataset exemplifies the necessity of multimodal understanding in categorizing online video content. The dataset is licensed under the Apache 2.0 License.
Given the substantial volume of the YouTube2M dataset, evaluation and fine-tuning on the entire dataset presents challenges such as runtime costs and catastrophic forgetting Huang et al. (2024); Zhai et al. (2024), where LLMs severely forget previously acquired information upon being trained on new data. To address potential biases and the predominance of YouTube data in tagging tasks, we strategically curated a subset of 2,000 examples from YouTube2M, aiming to mitigate any disproportionate influence of tagging tasks on the fine-tuning process. We partitioned the sampled dataset into training and test sets with an 80:20 ratio.
C.2 Misinformation datasets
@@ C.4 Memotion
Appendix D Details about Experiments
D.1 Implementation Details
Benchmark Evaluation For inference, we use Nucleus Sampling Holtzman et al. (2019) with a probability threshold of 0.9, a temperature of 1.0, and a maximum output length of 256 tokens. To account for the randomness in the generation process, we run each experiment with 3 random seeds and report the average results. All experiments were conducted on a server with 5 A100 80GB GPUs. The models are implemented using the Transformers library Wolf et al. (2020). We use the NLTK package Loper and Bird (2002) to calculate BLEU scores, the rouge [^4] https://github.com/pltrdy/rouge package to calculate ROUGE scores and the sentence-bert [^5] https://github.com/UKPLab/sentence-transformers package to calculate sentence embedding similarities, respectively.
D.2 Evaluation Metrics
@@
Task Acc Pre Rec F1
Misinformation 0.835 0.758 0.773 0.765
Humor 0.547 0.486 0.483 0.477
Sarcasm 0.717 0.544 0.519 0.503
Offensive 0.389 0.403 0.432 0.364
Sentiment 0.244 0.272 0.162 0.198
Hate Speech 0.670 0.626 0.695 0.614
Table 11: Results on GPT4V.
@@ D.3 Details on Models
Table 14: Accuracy of all models on the classification tasks. Best and 2nd best performances among the MLLMs are highlighted in bold and underline, respectively.